# The self-interaction effects on the Kerr black hole superradiance and their observational implications

### v1

@@ Abstract
Through the black hole (BH) superradiance, ultralight bosons can form dense clouds around rotating Kerr BHs. Certain ultralight bosons, such as axions and axion-like particles (promising dark matter candidates), naturally possess self-interactions, and thus may significantly modify the dynamics of the superradiance process. Previous studies on the detection or constraint of ultralight bosons through superradiance have largely neglected the self-interaction effects of bosons. In this work, we investigate the formation and evolution of self-interacting boson clouds in the full Kerr spacetime during BH superradiance. Using numerical methods, we compute the superradiant growth rate of boson clouds with self-interactions around Kerr BHs and quantitatively evaluate how the self-interaction strength of scalar bosons affects the growth rate. We also assess the evolution of the BH’s mass and spin. Our results reveal that, in addition to the superradiance-imposed upper bound on the boson cloud mass, self-interactions of ultralight bosons introduce a new, lower critical mass limit, beyond which the growth rate of the boson cloud approaches zero. This implies that the superradiance process terminates earlier when self-interactions are considered. Furthermore, we explore how self-interactions affect both the oscillation frequency of boson clouds in gravitational atoms and the frequency of gravitational wave (GW) emitted through cloud annihilation. The anticipated frequency shift could be detectable by the GW observatories. Given that self-interactions substantially alter the evolution of BH superradiance, their effects can significantly relax existing constraints on scalar boson models derived from superradiance. Taking the spin measurements from GW190412 and GW190517 as examples, we discuss the impact of self-interactions on constraint results in details.
@@ I Introduction
The ultralight bosons, such as the axions [1, 2, 3, 4, 5, 6, 7, 8] and axion-like particles [9], naturally correlate particle physics, cosmology, and gravity in many important processes, one of which is the superradiance mechanism [10, 11, 12, 13, 14, 15] of black hole (BH). These fundamental particles can serve as the well-motivated dark matter candidates [16, 17, 18, 19]. The superradiance process can amplify bosonic waves near a fast-spinning BH by extracting energy and angular momentum of BH and the macroscopic condensed boson cloud can form spontaneously. The boson cloud consisting of ultralight boson particles exhibits an exponential growth due to the boundary condition at the outer horizon of Kerr BH. The superradiant instability of ultralight bosons can be significant when the Compton wavelength of boson particle is comparable to the BH size, and consequently, the mass of boson cloud could reach up to $\sim 10\leavevmode\nobreak\ \textrm{\%}$ of the BH mass after the growth saturates. The black hole–boson cloud system, so-called gravitational atom, can form for either the isolated BH or the BH in compact binary.
@@
BH superradiance [10, 11, 12, 13, 14, 15], which minimally only requires the gravitational interaction of the ultralight bosons to induce, is a well-established probe in the search for the ultralight bosons and dark matter. The formation of the dense, macroscopic bosonic condensate consisting of ultralight particles around the BH can lead to various observable signatures. These effects caused by the superradiance include the modification on the dynamic of BH inspiral [20, 21, 22, 23, 24, 25, 26, 27], the continuous GW emission from the boson cloud [28, 29, 30, 31, 32, 33, 34, 35], and the spin-down of BH [36, 37, 38, 39, 40, 41, 42].
@@
The discovery of gravitational waves (GWs) [43] has initiated new perspectives to explore the dark matter through various GW experiments including LIGO [44], TianQin [45, 46, 47], LISA [48], and Taiji [49]. In order to investigate the signatures from the superradiant growth of ultralight bosons, it is important to rigorously understand the formation and dynamical evolution of the boson cloud, and a detailed evaluation of the black hole–boson cloud system is required to make the appropriate interpretation for the current and upcoming observations and the constraints on the boson parameter space from the experiments.
@@
Unlike in the purely gravitational scenario for single black hole–boson cloud system, the evolution of the gravitational atom can be influenced significantly due to the self-interactions of the scalar bosons [50, 51], other interactions with Standard Model particles [52, 53, 54, 55], and the tidal interaction from the macroscopic companion object [24, 25, 26, 56, 57, 58, 59, 60, 61]. Effects of self-interactions on the boson cloud evolution have attracted increasing attention [50, 51, 62] and become important in the experimental searches for ultralight bosons [63, 64, 41, 65, 66]. Ref. [50] addresses these questions with a detailed treatment of scalar self-interactions and considers the full dynamical evolution of a gravitational atom. Earlier than that, some of the studies argue during the growth of boson cloud, the strong attractive self-interactions would finally lead to the nonperturbative collapse of the cloud and consequently the drastic phenomenon called “bosenova” [36, 67, 68]. Nonetheless, the current study [50] proposes that when considering the energy exchange and level mixing between different bound states of the gravitational atom, the quartic self-interaction will cause the boson cloud to saturate at a new equilibrium with less occupation number. As a result, there is low probability for the bosenova to occur in the parameter space of interest. Overall, although analyzed at the semianalytic level, this study [50] demonstrates that the existence of self-interaction for bosons can dramatically alter the dynamical evolution of BH superradiance. The boson cloud can stay at a two-level equilibrium stage and the evolution of BH parameters is modified. In Ref. [50], the spectra of boson clouds in the gravitational atom are described by the hydrogenic modes in the non-relativistic approximation, and the results are widely valid in the regime where the spacetime curvature at the position of boson cloud can be neglected.
@@
In contrast to Ref. [50], Ref. [62] investigates the effects of the self-interactions on the saturated boson cloud in the nonlinear regime. In Ref. [62], the authors solve for the saturation stage of the gravitational atom for the dominant bound state considering the backreaction of the cloud on the metric, but ignore the effect of BH spin so as to simplify the spacetime to be spherically symmetric. The authors propose that, in comparison to the case without self-interaction, the mass of the self-interacting boson cloud in the saturated stage may increase by up to 70% when the coupling parameter lies within $-10^{-2}\lesssim\lambda\lesssim 10^{-2}$. Nonetheless, the additional backreaction effect of boson cloud on the metric is always negligible. Also, the authors find an approximate quadratic scaling relation between the mass of the boson cloud and the coupling parameter. However, for the ultralight bosons, if we consider a much smaller coupling parameter for which the energy scale of the decay constant is around the Grand Unified Theory scale or the Planck scale, the increment of the cloud mass due to the effects discussed in [62] would be insignificant. As in Ref. [50], we will focus on the case in which the energy scale of the decay constant for the ultralight bosons larger than $10^{13}\leavevmode\nobreak\ \mathrm{GeV}$ is considered in this work, since even a small value of self-interaction parameter can play an important role in the formation and evolution of boson cloud before it is saturated.
@@
We investigate the evolution of the gravitational atom from the initial stage that the boson cloud formation begins to the cease of cloud growth, taking into account the self-interaction effects on the dominant bound state of boson cloud. We solve the field equation of scalar bosons in the vicinity of BH and find the eigenfrequency for the dominated bound state in the full Kerr metric. We then analyze in detailed the growth rate of self-interacting boson cloud and the evolution of BH parameters. Our calculation is reliable for a relatively large decay constant of boson field and under the assumption that the growth rate of boson cloud is much slower than the oscillation frequency of boson field around the BH. In fact, we find that over a large range of parameter space of interest for the gravitational atom, these approximations are sufficiently valid. We calculate the results within a relativistic framework. Thus, the results can be applied to the gravitational atom with a large range of its parameters such as BH mass and boson mass. It is necessary to note that, if we adopt the non-relativistic approximation, the treatment of the self-interaction term for a single-state boson cloud in this study only leads to a quantitative shift in the energy potential of the bound state, as also analyzed in Ref. [50]. When we solve the boson field equation in the flat-spacetime limit, the deviation of the boson cloud growth rate disappears . Therefore, the modification on the growth rate of the boson cloud comes from not only the self-interactions of the bosons, but also the curvature nature of spacetime background.
@@
The paper is organized as follows. We first review the well-known superradiance process and introduce the self-interaction of scalar boson. We focus on the evaluating method to impose the quartic self-interaction term for scalar particles in Sec II. In Sec III, we apply the numerical approach to compute the superradiant growth rate of self-interacting boson clouds around the BH. Then, we evaluate how the self-interaction among bosons alter the GWs emitted by the boson clouds in Sec IV. In Sec V, we study in detail the potential impact of self-interactions on the existing observations and experiments, especially the relaxation of constraints on scalar bosons. We conclude and discuss future directions in Sec VI.
@@ II Superradiance and boson cloud with self-interaction
In this section, we present a concise overview of the superradiance phenomenology for a (free) scalar field and especially the quartic self-interaction term of the scalar field. We then focus on the numerical method to compute the spectra of a gravitational atom with considering the self-interaction of boson clouds. We adopt the units in which $\hbar=G=c=1$, unless otherwise specified.
@@ II.1 Black hole superradiance
The boundary condition at horizon of the Kerr BH leads to imaginary frequency of ultralight particles and then the exponential growth of the bound-state scalar bosons. In other words, a rapidly rotating BH can spontaneously transform energy and angular momentum to generate a macroscopic scalar cloud and form a gravitational atom through superradiance process [10, *1972SR, 12, 13, 14, 15]. In quantum field theory, the dynamics of a scalar boson field in the Kerr spacetime are governed by the Klein-Gordon equation. In general case, solving this equation numerically yields the eigenstates and the frequency eigenvalues of the scalar cloud. It has been demonstrated that the boson field exhibits a positive imaginary component in its frequency eigenvalue. This positive imaginary part results in an exponentially growing solution for the boson cloud when the Compton wavelength of the boson is comparable to the scale of the BH. Nevertheless, up to date, the numerical studies on Klein-Gordon solution of boson cloud’s formation on the Kerr background has not taken into account the self-interaction of the field.
[…]
$\dfrac{\alpha}{m}<\dfrac{\chi}{2\left(1+\sqrt{1-\chi^{2}}\right)}\,\,,$ (1)
where $\chi$ is the dimensionless spin of rotating BH ($10<\chi<$ 1) and $m$ is the magnetic quantum number of the scalar cloud.
@@
The superradiance rates—–imaginary parts of bound state frequencies—– depend on boson mass and BH spin. Specifically, for $2p$ level bound state ($\,\left|211\right\rangle$ state), the boson cloud grows and saturates within the timescale of $\tau_{\text{SR}}=1\leavevmode\nobreak\ \text{day}\ (\frac{M_{\text{BH}}}{100\leavevmode\nobreak\ {\rm M_{\odot}}})(\frac{{0.2}}{{}_{\,}{\displaystyle\alpha}_{\,}})^{9}(\frac{0.9}{{\chi}})$ [31]. When exponential growth stops the scalar cloud extracts part of angular momentum from BH, and the number of scalar boson could be evaluated by [29]
$N_{a\rm{}}\sim 10^{78}\left(\dfrac{\varDelta\chi}{0.1}\right)\dfrac{1}{m}\left(\dfrac{M_{\text{BH}}}{100\leavevmode\nobreak\ {\rm M_{\odot}}}\right)^{2}\,\,,$ (2)
with $\varDelta\chi$ the spin difference.
@@
When ${\alpha}/{m}$ is much smaller than unity, the angular frequency eigenvalue of bound state can be solved perturbatively. If assuming the boson cloud to be non-relativistic, we can factor out the phase term in the field function $\phi$ related to its mass, and the leaf scalar function satisfies the Schrödinger-like equation. By solving the equation, one can in principle determine the eigenstate and (real) frequency eigenvalue of scalar bosons analytically. The eigenstates of boson clouds are analogous to hydrogenic bound states in the non-relativistic limit. On the other hand, in the limit that $\alpha$ is much smaller than unity, one can employ the Detweiler’s approximation in order to derive the imaginary frequency of boson clouds, and give the analytical expressions of the growth rates [69, 70].
@@ II.2 Numerical solution for the ultralight scalar bosons
For the scalar field, the radial part and the angular part of the field function $R(r)$ and $S(\theta)$ satisfy the confluent Heun equation and the spheroidal harmonic equation,
@@
where the first several coefficients $u_{n}$ had been calculated by Seidel [71]. On the Kerr background, due to the behaviour of the scalar field at the horizon and infinity, we can expand the radial function $R(r)$ as

### v2

# Self-interaction effects on the Kerr black hole superradiance and their observational implications

@@ Abstract
Through the black hole (BH) superradiance, ultralight bosons can form dense clouds around rotating Kerr BHs. Certain ultralight bosons, such as axions and axion-like particles (promising dark matter candidates), naturally possess self-interactions, and thus may significantly modify the dynamics of the superradiance process. Previous studies on the detection or constraint of ultralight bosons through superradiance have usually neglected the self-interaction effects of bosons. In this work, we investigate the formation and evolution of self-interacting boson clouds in the full Kerr spacetime during BH superradiance. Using numerical methods, we compute the superradiant growth rate of boson clouds with self-interactions around Kerr BHs and quantitatively evaluate how the self-interaction strength of scalar bosons affects the growth rate. We also assess the evolution of the BH’s mass and spin. Our results reveal that, in addition to the superradiance-imposed upper bound on the boson cloud mass, self-interaction of ultralight bosons introduces a new, lower critical mass limit, beyond which the growth rate of the boson cloud approaches zero. This implies that the superradiance process terminates earlier when self-interaction is considered. Furthermore, we explore how self-interaction affects both the oscillation frequency of boson clouds in gravitational atoms and the frequency of gravitational wave (GW) emitted through cloud annihilation. The anticipated frequency shift might be detectable by the GW observatories. Given that self-interaction substantially alters the evolution of BH superradiance, its effects can significantly relax existing constraints on scalar bosons derived from superradiance. Taking the spin measurements from GW190412 and GW190517 as examples, we discuss the impact of self-interaction on constraint results in details.
@@ I Introduction
The ultralight bosons, such as the axions [1, 2, 3, 4, 5, 6, 7, 8] and axion-like particles [9], naturally correlate particle physics, cosmology, and gravity in many important processes, one of which is the superradiance mechanism [10, 11, 12, 13, 14, 15, 16] of Kerr black hole (BH). These fundamental particles can serve as the well-motivated dark matter candidates [17, 18, 19, 20]. The superradiance process can amplify bosonic waves near a fast-spinning BH by extracting energy and angular momentum of BH and the macroscopic condensed boson cloud can form spontaneously. The boson cloud consisting of ultralight boson particles exhibits an exponential growth due to the boundary condition at the outer horizon of Kerr BH. The superradiant instability of ultralight bosons can be significant when the Compton wavelength of the boson particle is comparable to the BH size, and consequently, the mass of boson cloud could reach up to $\sim 10\textrm{\%}$ of the BH mass after the growth saturates. The BH–boson cloud system, so-called gravitational atom, can form for either the isolated BH or the BH in compact binary.
@@
BH superradiance [10, 11, 12, 13, 14, 15, 16], which minimally only requires the gravitational interaction of the ultralight bosons to be induced, is a well-established probe in the search for the ultralight bosons and dark matter. The formation of the dense, macroscopic bosonic condensate consisting of ultralight particles around the BH can lead to various observable signatures. These effects caused by the superradiance include the modification on the dynamic of BH inspiral [21, 22, 23, 24, 25, 26, 27, 28], the continuous GW emission from the boson cloud [29, 30, 31, 32, 33, 34, 35, 36], and the spin-down of BH [37, 38, 39, 40, 41, 42, 43].
@@
The discovery of gravitational waves (GWs) [44] has initiated new perspectives to explore the dark matter through various GW experiments including LIGO [45], TianQin [46, 47, 48], LISA [49], and Taiji [50]. In order to investigate the signatures from the superradiant growth of ultralight bosons, it is important to rigorously understand the formation and dynamical evolution of the boson cloud, and a detailed evaluation of the BH–boson cloud system is required to make the appropriate interpretation for the current and upcoming observations and the constraints on the boson parameter space from the experiments.
@@
Unlike in the purely gravitational scenario for single BH–boson cloud system, the evolution of the gravitational atom can be influenced significantly due to the self-interactions of the scalar bosons [51, 52], other interactions with Standard Model particles [53, 54, 55, 56, 57, 58], and the tidal interaction from the macroscopic companion object [25, 26, 27, 59, 60, 61, 62, 63, 64]. Effects of self-interactions on the boson cloud evolution have attracted increasing attention [51, 52, 65, 66, 67] and become important in the experimental searches for ultralight bosons [68, 69, 42, 70, 71]. Ref. [51] addresses these questions with a detailed treatment of scalar self-interactions and considers the full dynamical evolution of a gravitational atom. Earlier than that, some of the studies argue during the growth of boson cloud, the strong attractive self-interactions would finally lead to the nonperturbative collapse of the cloud and consequently the drastic phenomenon called “bosenova” [37, 72, 73]. Nonetheless, the current study [51] proposes that when considering the energy exchange and level mixing between different bound states of the gravitational atom, the quartic self-interaction will cause the boson cloud to saturate at a new equilibrium with less occupation number. As a result, there is low probability for the bosenova to occur in the parameter space of interest. Overall, although analyzed at the semi-analytic level, this study [51] demonstrates that the existence of self-interaction for bosons can dramatically alter the dynamical evolution of BH superradiance. The boson cloud can stay at a two-level equilibrium stage and the evolution of BH parameters is modified. In Ref. [51], the spectra of boson clouds in the gravitational atom are described by the hydrogenic modes in the non-relativistic approximation, and the results are widely valid in the regime where the spacetime curvature at the position of boson cloud can be neglected.
@@
In contrast to Ref. [51], Ref. [65] investigates the effects of the self-interactions on the saturated boson cloud in the nonlinear regime. In Ref. [65], the authors solve for the saturation stage of the gravitational atom for the dominant bound state considering the backreaction of the cloud on the metric, but ignore the effect of BH spin so as to simplify the spacetime to be spherically symmetric. The authors propose that, in comparison to the case without self-interaction, the mass of the self-interacting boson cloud in the saturated stage may increase by up to 70% when the coupling parameter lies within $10^{-2}\lesssim\lambda\lesssim 10^{-2}$. Nonetheless, the additional backreaction effect of boson cloud on the metric is always negligible. Also, the authors find an approximate quadratic scaling relation between the mass of the boson cloud and the coupling parameter. However, for the ultralight bosons, if we consider a much smaller coupling parameter for which the energy scale of the decay constant is around the Grand Unified Theory scale or the Planck scale, the increment of the cloud mass due to the effects discussed in [65] would be insignificant. As in Ref. [51], we will focus on the case in which the energy scale of the decay constant for the ultralight bosons larger than $10^{13}\penalty 10000\ \mathrm{GeV}$ is considered in this work, since even a small value of self-interaction parameter can play an important role in the formation and evolution of boson cloud before it is saturated.
@@
We investigate the evolution of the gravitational atom from the initial stage that the boson cloud formation begins to the cease of cloud growth, taking into account the self-interaction effects on the dominant bound state of boson cloud. We solve the field equation of scalar bosons in the vicinity of BH and find the eigenfrequency for the dominated bound state on the full Kerr metric. We then analyze in detailed the growth rate of self-interacting boson cloud and the evolution of BH parameters. Our calculation is reliable for a relatively large decay constant of boson field and under the assumption that the growth rate of boson cloud is much slower than the oscillation frequency of boson field around the BH. In fact, we find that over a large range of parameter space of interest for the gravitational atom, these approximations are sufficiently valid. We calculate the results within a relativistic framework. Thus, the results can be applied to the gravitational atom with a large range of its parameters such as BH mass and boson mass. It is necessary to note that, if we adopt the non-relativistic approximation, the treatment of the self-interaction term for a single-state boson cloud in this study only leads to a quantitative shift in the energy potential of the bound state, as also analyzed in Ref. [51]. When we solve the boson field equation in the flat-spacetime limit, the deviation of the boson cloud growth rate disappears . Therefore, the modification on the growth rate of the boson cloud comes from not only the self-interactions of the bosons, but also the curvature nature of spacetime background.
@@
The paper is organized as follows. We first review the well-known superradiance process and introduce the self-interaction of scalar boson. We focus on the evaluating method to impose the quartic self-interaction term for scalar particles in Sec II. In Sec III, we apply the numerical approach to compute the superradiant growth rate of self-interacting boson clouds around the BH. Then, we evaluate how the self-interaction among bosons alters the GWs emitted by the boson clouds in Sec IV. In Sec V, we study in detail the potential impact of self-interactions on the existing observations and experiments, especially the relaxation of constraints on scalar bosons. We conclude and discuss future directions in Sec. VI.
@@ II Superradiance and boson cloud with self-interaction
In this section, we present a concise overview of the superradiance phenomenology for a (free) scalar field and especially the quartic self-interaction term of the scalar field. We then focus on the numerical method to compute the spectra of a gravitational atom considering the self-interaction of boson clouds. We adopt the units in which $\hbar=G=c=1$, unless otherwise specified.
@@ II.1 Black hole superradiance
The boundary condition at horizon of the Kerr BH leads to imaginary frequency of ultralight particles and then the exponential growth of the bound-state scalar bosons. In other words, a rapidly rotating BH can spontaneously transform energy and angular momentum to generate a macroscopic scalar cloud and form a gravitational atom through superradiance process [10, *1972SR, 12, 13, 14, 15, 16]. In quantum field theory, the dynamics of a scalar boson field in the Kerr spacetime are governed by the Klein-Gordon equation. In general case, solving this equation numerically yields the eigenstates and the frequency eigenvalues of the scalar cloud. It has been demonstrated that the boson field exhibits a positive imaginary component in its frequency eigenvalue. This positive imaginary part results in an exponentially growing solution for the boson cloud when the Compton wavelength of the boson is comparable to the scale of the BH. Nevertheless, up to date, the numerical studies on Klein-Gordon solution of boson cloud’s formation on the Kerr background has not taken into account the self-interaction of the field.
$V_{\rm free}(\phi)=\frac{1}{2}{m_{a}^{2}}\phi^{2}\,\,.$ (2)
Therefore one can have the Klein-Gordon equation
$(\nabla^{2}-m_{a}^{2})\phi=0$ (3)
for the free field. The metric convention here is (,+,+,+)(-,+,+,+).
In the gravitational atom, the gravitational fine-structure constant is defined as $\alpha={M_{\text{BH}}}m_{a}$, where $m_{a}$ is the mass of scalar boson and $M_{\text{BH}}$ is the Kerr BH mass. This parameter plays a critical role in determining the efficiency of the superradiant instability and the time scale of scalar cloud formation. For the scalar field in the vicinity of the Kerr BH, the boundary conditions at horizon and infinity lead to imaginary eigenfrequency of the bosons. The boson cloud with real eigenfrequency $\omega_{\text{Re}}$ and magnetic quantum number $m$ could extract energy and angular momentum from BH spontaneously, if the initial spin of BH is greater than the critical value that triggers the superradiance process with the following superradiance condition [15, 30, 16] reads
$\omega{}{}_{\text{Re}}<\dfrac{m\chi}{2r_{+}}\,\,$, (4)
where $\chi$ is the dimensionless spin of rotating BH ($0<\chi<1$) and $r_{+}={M_{\text{BH}}}\left(1+\sqrt{1-\chi^{2}}\right)$ is the radius of the outer event horizon. Although in some previous references this condition was derived analytically, it could also be demonstrated by numerically solving the superradiant instability rate, see Fig. 1 for details. Since a significant cloud growth rate requires $\alpha\sim\mathcal{O}(10^{-2})\textrm{--}\mathcal{O}(1)$, for the mass range of the astrophysical BHs, we are able to investigate the boson mass ranging from $10^{-21}$ to $10^{-10}\penalty 10000\ \text{eV}$. In this work, for simplicity, we choose the benchmark BH mass $M_{\text{BH}}=25,\ 40,\ 60,\text{and }100\penalty 10000\ {\rm M_{\odot}}$, and focus on the boson mass range $10^{-14}\,\textrm{--}\,10^{-12}\penalty 10000\ \text{eV}$.
@@
The superradiance rates—–imaginary parts of bound state frequencies—–depend on boson mass and BH spin. Specifically, for $2p$ level bound state ($\left|211\right\rangle$ state), the boson cloud grows and saturates with the timescale of $\tau_{\text{SR}}=1\penalty 10000\ \text{day}\left(\frac{M_{\text{BH}}}{40\penalty 10000\ {\rm M_{\odot}}}\right)\left(\frac{{0.2}}{{}_{\,}{\alpha}_{\,}}\right)^{9}\left(\frac{0.9}{{\chi}}\right)$ [32]. When exponential growth stops the scalar cloud extracts part of angular momentum from BH, and the occupation number of scalar boson could be evaluated by [30]
$N_{a\rm{}}\sim 10^{77}\left(\dfrac{\varDelta\chi}{0.1}\right)\dfrac{1}{m}\left(\dfrac{M_{\text{BH}}}{40\penalty 10000\ {\rm M_{\odot}}}\right)^{2}\,\,$, (5)
with $\varDelta\chi$ the spin difference of the BH from the initial stage when the boson cloud formation begins. Boson cloud formation depletes BH spin through continuous transfer of angular momentum from BH to the bound state. Eq. (5) demonstrates the conservation of angular momentum.
@@
When ${\alpha}/{m}$ is much smaller than unity, the angular frequency eigenvalue of bound state can be solved perturbatively. If assuming the boson cloud to be non-relativistic, we can factor out the phase term in the field function $\phi$ related to its mass, and the remaining scalar function satisfies the Schrödinger-like equation. By solving the equation, one can in principle determine the eigenstate and (real) frequency eigenvalue of scalar bosons analytically. The eigenstates of boson clouds are analogous to hydrogenic bound states in the non-relativistic limit. On the other hand, in the limit that $\alpha$ is much smaller than unity, one can employ the Detweiler’s approximation in order to derive the imaginary frequency of boson clouds, and give the analytical expressions of the growth rates [74, 75].
@@ II.2 Numerical solution for the ultralight scalar bosons
For the scalar field, the radial and angular part of the field function $R(r)$ and $S(\theta)$ satisfy the confluent Heun equation and the spheroidal harmonic equation,
@@
where the first several coefficients $u_{n}$ had been calculated by Seidel [76]. On the Kerr background, due to the behavior of the scalar field at the horizon and infinity, we can expand the radial function $R(r)$ as